OpenStudy (anonymous):

A firecracker shoots up from a hill 140 feet high with an initial speed of 100 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? Answer 5 seconds 7 seconds 9 seconds 11 seconds HELP D;

jimthompson5910 (jim_thompson5910):

H(t) = -16t^2 + vt + s H(t) = -16t^2 + 100t + 140 0 = -16t^2 + 100t + 140 -16t^2 + 100t + 140 = 0 -4(4t^2 - 25t - 35) = 0 -4(4t^2 - 25t - 35) = 0 4t^2 - 25t - 35 = 0/(-4) 4t^2 - 25t - 35 = 0 Now use the quadratic formula to solve for t t = (-b+-sqrt(b^2-4ac))/(2a) t = (-(-25)+-sqrt((-25)^2-4(4)(-35)))/(2(4)) t = (25+-sqrt(625-(-560)))/(8) t = (25+-sqrt(1185))/8 t = (25+sqrt(1185))/8 or t = (25-sqrt(1185))/8 t = 7.42797 or t = -1.17797 So it will take roughly 7 seconds for the firecracker to hit the ground.